### v1

@@ 2.1 The Problem
To make this problem more accessible, especially for students, we modified it slightly to make it easier to find any feasible solution at all, and to have some more tractable but non-trivial solutions available, and, thus, a more continuous difficulty spectrum. First, we changed the problem to accept obtuse triangles, splitting the objective into first minimizing the number of obtuse triangles, and then minimizing the number of Steiner points. This makes any triangulation feasible and serves as a starting point for iterative improvement. Second, we introduced a boundary to the triangulation area, which encapsulate the area of interest into a simple polygon. While this may first sound like a complicating constraint, it actually simplifies the problem, as Steiner points placed on the boundary are less of a risk to create further obtuse triangles that propagate through the whole instance.
@@ 2.4 Instances
The generation of appropriate instances is critical for any challenge. Instances that are too easily solved undermine the complexity of the problem, rendering it trivial. Conversely, instances that require significant computational resources for preprocessing or for finding viable solutions may offer an unfair advantage to teams with superior computing facilities. This issue is amplified when the instance set is so large that it exceeds the management capacity of a limited number of computers.
@@ 3 Outcomes
    Team Naughty NOTers: Mikkel Abrahamsen, Florestan Brunck, Jacobus Conradi, Benedikt Kolb and André Nusser [1].

### v2

@@ 2.1 The Problem
To make this problem more accessible, especially for students, we modified it slightly to make it easier to find any feasible solution at all, and to have some more tractable but non-trivial solutions available, and, thus, a more continuous difficulty spectrum. First, we changed the problem to accept obtuse triangles, splitting the objective into first minimizing the number of obtuse triangles, and then minimizing the number of Steiner points. This makes any triangulation feasible and serves as a starting point for iterative improvement. Second, we introduced a boundary to the triangulation area, which encapsulates the area of interest into a simple polygon. While this may first sound like a complicating constraint, it actually simplifies the problem, as Steiner points placed on the boundary are less of a risk to create further obtuse triangles that propagate through the whole instance.
@@ 2.4 Instances
The generation of appropriate instances is critical for any challenge. Instances that are too easily solved undermine the complexity of the problem, rendering it trivial. Conversely, instances that require significant computational resources for preprocessing or for finding viable solutions may offer an unfair advantage to teams with superior computing facilities. This issue is amplified when the instance set is so large that it exceeds the management capacity when using a limited number of computers.
@@ 3 Outcomes
    Team Naughty NOTers: Mikkel Abrahamsen, Florestan Brunck, Jacobus Conradi, Benedikt Kolbe and André Nusser [1].